OpenStudy (anonymous):

Conics: Write in standard form 4x^2+4y^2+32x-18y+37=0

myininaya (myininaya):

its a circle! so we need to put in the form (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Complete the square for x and y.

OpenStudy (anonymous):

For completing the square I was confused because of the 7 you get when you factor the 4 out for y

myininaya (myininaya):

divide everything by 4 \[x^2+y^2+8x-\frac{9}{2}y+\frac{37}{4}=0\]

myininaya (myininaya):

\[x^2+8x+(\frac{8}{2})^2+y^2-\frac{9}{2}y+(\frac{9}{2*2})^2+\frac{37}{4}=16+\frac{81}{16}\]

myininaya (myininaya):

\[(x+4)^2+(y-\frac{9}{4})^2=\frac{337}{16}\]

OpenStudy (anonymous):

So if I am also looking for the center and radius it would be -4, \[(9 \ 4)\] and the radius would be \[\sqrt{337/16}\] ?

OpenStudy (anonymous):

*-4, (9/4)

myininaya (myininaya):

yep

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

Thank you so much